OpenStudy (anonymous):

help plz i need to evaluate ((2^2x^2)/(xy^2))^2 for x=3 and y=2

OpenStudy (mathmale):

\[((2^2x^2)/(xy^2))^2\] is how your algebraic expression appears in Equation Editor. You have done a good job of presenting this problem by using parentheses liberally. Parentheses are of great importance here because they indicate which parts of the expression to evaluate first. First of all, you have outermost parentheses with a quotient inside. Order of operations rules require that you evaluate what's inside those outermost parentheses first (although in this particular problem you could do the work in a different order). Assuming that we evaluate what's inside the outermost parentheses first: 1) Evaluate 2^2*x^2 for x=3. Could you do that now, please?

OpenStudy (mathmale):

Welcome to OpenStudy! Bet you're busy exploring problems that other students have posted. When you return to your own problem here, evaluate\[2^2x^2\] for x=3, and then (2) evaluate \[xy^2\] for x=3, y=2. What do you think the next step should be?

OpenStudy (mathmale):

I see you are now offline. Respond when you come back online, if you wish. Welcome to OpenStudy.

OpenStudy (anonymous):

i dont know

OpenStudy (mathmale):

I am on a phone call at the moment and will return to this conversation a bit later. What is 2^2? What is 3^2? I'm sure you've seen problems like these before.

OpenStudy (mathmale):

To help you get started, 2^2 = 4, or, written using Equation Editor,\[2^2=4\] Now please evaluate 3^2. That's "3 squared."

OpenStudy (mathmale):

sunrise, you can get a lot of help here on OpenStudy, but you need to be personally involved and not hope that someone will explain the solution of a given problem from start to finish. 2^2 means 2 times 2, and the correct result / answer is 4. 3^2 means 3 times 3. What is the correct result?

OpenStudy (anonymous):

9

OpenStudy (mathmale):

Good. What is (2^2)(3^2)?

OpenStudy (anonymous):

36

OpenStudy (mathmale):

Good. So you now have the value of (2^2*3^2), the numerator, that's inside the outermost set of parentheses. Now look at the denominator. You have xy^2. Translate that into words, please.

OpenStudy (anonymous):

3 * 2 to the second power

OpenStudy (mathmale):

"xy^2" or \[xy^2\]

OpenStudy (mathmale):

is the product of x times the square of y. "Order of operations" rules require that you evaluate the exponential (y^2) BEFORE you multiply by x. If y = 2, what is y^2? If x = 3, what is (x) (y^2)?

OpenStudy (anonymous):

y^2 is 4

OpenStudy (mathmale):

Right. and what is the value of x(4), if x=3?

OpenStudy (anonymous):

12

OpenStudy (mathmale):

Very good. So now you have\[(\frac{ 36 }{ 12 })^2\]First, reduce the quantity in parentheses. Then raise the result to the power 2.

OpenStudy (mathmale):

Be certain that you know where that 36 and that 12 came from.

OpenStudy (mathmale):

OpenStudy tells me that you are looking at a different math problem now, not on the one that we were discussing. Partly because of that, I'm logging off OpenStudy for the time being. But I'll be back on later. Best to stick with ONE problem at a time if you want people to continue helping you. See you later.

OpenStudy (anonymous):

im on